### v1

# From Specificity to Generality: Revisiting Generalizable Artifacts in Detecting Face Deepfakes

Abstract
Corresponding Author.
Detecting deepfakes has been an increasingly important topic, especially given the rapid development of AI generation techniques. In this paper, we ask: How can we build a universal detection framework that is effective for most facial deepfakes? One significant challenge is the wide variety of deepfake generators available, resulting in varying forgery artifacts (e.g., lighting inconsistency, color mismatch, etc). But should we “teach” the detector to learn all these artifacts separately? It is impossible and impractical to elaborate on them all. So the core idea is to pinpoint the more common and general artifacts across different deepfakes. Accordingly, we categorize deepfake artifacts into two distinct yet complementary types: Face Inconsistency Artifacts (FIA) and Up-Sampling Artifacts (USA). FIA arise from the challenge of generating all intricate details, inevitably causing inconsistencies between the complex facial features and relatively uniform surrounding areas. USA, on the other hand, are the inevitable traces left by the generator’s decoder during the up-sampling process. This categorization stems from the observation that all existing deepfakes typically exhibit one or both of these artifacts. To achieve this, we propose a new data-level pseudo-fake creation framework that constructs fake samples with only the FIA and USA, without introducing extra less-general artifacts. Specifically, we employ a super-resolution to simulate the USA, while design a Blender module that uses image-level self-blending on diverse facial regions to create the FIA. We surprisingly found that, with this intuitive design, a standard image classifier trained only with our pseudo-fake data can non-trivially generalize well to unseen deepfakes.
1 Introduction
Figure 1: Illustration of the two identified general forgery artifacts across different face deepfakes: (a) Face Inconsistency Artifacts and (b) Up-Sampling Artifacts. We show that the existing face deepfakes typically exhibit both or one of these two artifacts.
Methods Alteration artifact Introduced Artifacts
global local Face Inconsistency Up-Sampling
face-level region-level
Face Xray [44], PCL [93], OST [4]
SLADD [3]
SBI [67]
SeeABLE [41]
RBI  [68]
Ours
Table 1: Comparison of our framework and Sota methods using pseudo-deepfake synthesis. Our method differs from previous methods in terms of the level at which alterations are applied (local vs global), the introduced artifacts, and the level at which facial inconsistencies are applied.
In recent years, A growing number of facial manipulation techniques have advanced due to the rapid development of generative models [39, 40, 62, 80], which facilitate the production of highly realistic and virtually undetectable alterations. As a result, the risks of personal privacy being violated and the spread of misinformation have increased significantly. Therefore, developing a universal deepfake detector that can be used to detect the existing diverse fake methods has become an urgent priority.
In this paper, we pose the question: How can we build a universal detection framework that is effective for most facial deepfakes? A significant challenge lies in the wide range of deepfake generators, which lead to different types of forgery artifacts such as lighting inconsistency and color mismatch. Specifically, a large number of existing works are dedicated to creating detection algorithms that are carefully designed to identify counterfeit artifacts within specific handcrafted designs, including eye blinking [46], pupil morphology [26], and corneal specularity [33]. However, these methods are mainly effective for identifying particular forgery techniques. But should we train the detector to learn each of these artifacts separately? It is impossible and impractical to cover them all. So, where should we start from?
The core idea is to encourage the detection model to learn the more generalizable artifacts. As indicated in previous works [84, 67, 75], different fakes might share some common forgery patterns, and the unlimited deepfake artifacts could be generalized and summarized by the limited generalizable artifacts. Motivated by this, we conduct an in-depth preliminary exploration of the common patterns of existing deepfake artifacts. After our preliminary exploration, we identify two generalizable deepfake artifacts and categorize them into two distinct yet complementary categories: Face Inconsistency Artifacts (FIA) and Up-Sampling Artifacts (USA), as depicted in Figure 1. FIA represents an inherent limitation that deepfakes struggle to overcome, originating from the generator’s incapacity to precisely reproduce facial attributes and nuances, as well as the inescapable disparities between regions induced by amalgamation processes. In contrast to FIA, USA is not readily apparent to the unaided eye and arises from an up-sampling process that the generator’s decoder cannot adequately substitute. Considering the intrinsic necessity for decoding and up-sampling within the generator’s operations, every facial manipulation method invariably introduces Up-Sampling Artifacts (USA). Numerous prior investigations have substantiated this observation [22, 70]. To date, all existing deepfakes characteristically display one or both of these artifact categories.
To enable the model to simultaneously capture FIA and USA, we propose a sophisticated data augmentation method: FIA-USA, which leverages super-resolution models [76] and autoencoders [61] to introduce Up-Sampling Artifacts (USA) by reconstructing the face, and employs a strategy of generating multiple masks to blend the reconstructed face with the original, thereby introducing Face Inconsistency Artifacts (FIA) by creating discrepancies at both the facial and regional levels. Augmented with the aforementioned dual artifacts, our method effectively conditions the model to generalize across unseen deepfakes. Furthermore, to maximize the efficacy of FIA-USA, we introduce two complementary technique: Automatic Forgery-aware Feature Selection (AFFS), which streamlines feature dimensionality and augments feature discernment through adaptive selection. Region-aware Contrastive Regularization (RCR), by juxtaposing the features of the manipulated region against those of the authentic region, RCR enhances the model’s focus on the upsampling traces within the manipulated areas and the inconsistencies between different regions.
Our method surpasses the most recent state-of-the-art data augmentation methods [44, 67, 79, 86, 68, 43] on traditional deepfake benchmarks, and we have realized substantial improvements in generation deepfake benchmarks when compared to the baseline method. In addition, we perform cross-manipulation and cross-dataset detection evaluation on unseen data to align with real-world detection scenarios. Our method has shown enhanced generalization and robustness in the context of cross dataset generalization.
2 Related Work
Deepfake Generation Methods.
In the field of forgery detection, the broad definition of deepfake detection encompasses both forged face image / video detection and generated image / video detection. In this paper, deepfake detection specifically refers to the detection of fake faces. Based on the methods by which fake faces are generated, we categorize deepfakes into two types: traditional deepfake, and the generative deepfake. Traditional deepfake methods typically involve face replacement [47, 13, 18] and face reenactment [72, 73]. Facial replacement typically involves identity exchange using autoencoder-based [13] or graph based exchange methods [18], while facial reenactment employs reconstruction techniques to exchange faces from the source video to the target video while maintaining consistency in the identity of the target person. Unlike traditional deepfakes that involve local tampering, generative deepfakes directly generate fake faces through generative models [38, 59, 61, 37]. As discussed in some works [87, 86], the detection methods using RGB data augmentation methods have poor performance in this type of forgery, as these generation methods do not involve mixing operations. More importantly, an increasing number of facial tampering methods [40, 39, 40, 80] are evolving based on generation models, which implies that the fake faces generated by these new tampering methods will contain both the artifacts introduced by the generative models and the blend boundaries brought by the blend operations. To date, Most of the work only focuses on one forgery method and is ineffective for other forgery methods.
Classic Deepfake Detection.
Classic deepfake detection can be categorized into multiple aspects: On the one hand, deepfake image detection, which can also be applied to deepfake video detection by computing the average value of multiple frames. Typical methods encompass data augmentation [44, 67, 86, 93], frequency-domain analysis [60, 36, 53, 24], identity information [17, 9, 74], contrastive learning [23, 91], network structure improvement [92, 17], watermark [89, 77], robustness [42, 55], interpretability [16, 66] and so forth. On the other hand, deepfake video detection innovative video representation formats [83], video-level data augmentation [79], self-supervised learning [2, 29, 96], anomaly detection [19, 20], biological signal [81, 8], multi-modal based [58]. In addition to the above two aspects, the the remainder of the work involves forgery localization [55, 27], multi task learning [55, 90], real scenario [69],adversarial attack [48], and so on.
3 Method
In this section, we begin by outlining two distinct and prevalent types of forgery artifacts found in deepfakes: Face Inconsistency Artifacts (FIA) and Up-Sampling Artifacts (USA). We then detail the construction of data augmentation techniques (FIA-USA) designed to enable the model to learn to detect these artifacts concurrently. Subsequently, we present the Automatic Forgery-aware Feature Selection (AFFS) and Region-aware Contrastive Regularization (RCR).
3.1 Face Inconsistency Artifacts (FIA)
As depicted in Figure 1, Face Inconsistency Artifacts (FIA) frequently occur in deepfakes due to the inherent difficulty in rendering intricate details, which often results in discrepancies between intricate facial features and the more uniform surrounding regions. In essence, until deepfake technology advances to the point where it can flawlessly replicate real faces, these artifacts are likely to persist. We will now explore the origins of Face Inconsistency Artifacts by categorizing deepfake techniques into two domains: Full Face Synthesis and Face Swapping. In the case of full face synthesis deepfakes, the presence of Face Inconsistency Artifacts is primarily contingent upon the capabilities of the generator, which, regardless of their magnitude, are invariably present. For face swapping, these artifacts arise from two primary sources: 1). the efficacy of the generator. 2). the blending seam created during the face swapping’s blending procedure. Furthermore, we can categorize face swapping into two main types based on the forged area: macro-editing and micro-editing. Macro-editing can be further classified into two subtypes: editing based on 4 facial keypoints and editing based on 81 keypoints, whereas micro-editing allows for precise adjustments to each facial feature. In section B of the appendix, we provide a detailed description of the classification of forgery techniques and their differences.
3.2 Up-Sampling Artifacts (USA)
Conversely, the artifacts invariably present in the forged regions produced by the forgery generator are due to the upsampling process within the generator’s decoder, known as Up-Sampling Artifacts (USA). The extent of USA introduced can differ among various generators. To investigate more universal upsampling artifacts, we categorize forgery techniques into two types: One category generates images of forged regions that are as clear and free of blur as the original image (using Diffusion models [61] and GANs [38] as the main methods, and upsampling the latent code through autoencoder (AE) to generate images). And another type result in blurred tampered areas, requiring additional facial super-resolution models (SR) to process the blurred areas.
3.3 FIA-USA
We propose a new data augmentation method to generate pseudo-fake training samples, aiming to introduce FIA and USA present in forged faces. The core idea is to introduce FIA and USA simultaneously by blending the real image $I_{R}$ and the reconstructed image $I_{F}$, as follows:
$I_{PF}=M\odot I_{F}+(1-M)\odot I_{R},$ (1)
where $\odot$ represents the element wise multiplication operator, $M$ represents the mask of the manipulated area, which varies between 0 and 1. We considered two ways of generating masks to cover all possible occurrences of FIA: 1) Mask erosion on the convex hull of the facial 81 or 4 keypoints, aimed at simulating macro-editing; 2) Random combination of facial features, aimed at simulating micro-editing. Consequently, we address generic USA through three approaches: 1. Employ an autoencoder (AE) to reconstruct the entire image [61]. 2. Enhance the facial details using a super-resolution model [76] (SR). 3. Do not reconstruct the original image. Formally speaking, we introduced different levels of FIA through blending operations and reconstructed images, while introducing USA for forged areas through reconstructed images. Please refer to  Algorithm 1 for the complete algorithmic process. Table 1 compares the differences between our proposed FIA-USA method and the previous blend-based data augmentation method. In section C of the appendix, we provide more information on blend-based data augmentation and FIA-USA.
Algorithm 1 Pseudocode for FIA-USA
Input: Base image $I$ of size $(H,W,3)$, facial landmarks $L$ of size $(81,2)$
Output: Image $I$ with FIA and USA
1:def $\mathcal{T}$($I$) : $\triangleright$ Source-Target Augmentation
2:     $I\leftarrow\text{ColorTransform}(I)$
3:     $I\leftarrow\text{FrequencyTransform}(I)$
4:     return $I$
5:def $Recon$($I$) : $\triangleright$ Reconstruct the source image
6:     if $\text{Uniform}(\text{min}=0,\text{max}=1)\in[0,0.3]$ :
7:         $I_{s}\leftarrow AE(I)$
8:     else if $\text{Uniform}(\text{min}=0,\text{max}=1)\in(0.3,0.5]$ :
9:         $I_{s}\leftarrow SR(I)$
10:     else
11:         $I_{s}\leftarrow I$
12:     return $I$
13:$I_{s},I_{t}\leftarrow Recon(I),I$
14:if $\text{Uniform}(\text{min}=0,\text{max}=1)<0.5$ :
15:     $I_{s},I_{t}\leftarrow\mathcal{T}(I_{s}),I_{t}$
16:else :
17:     $I_{s},I_{t}\leftarrow I_{s},\mathcal{T}(I_{t})$
18:$M\leftarrow\text{CombineFacialFeatures}(L)$ or $\text{ConvexHull}(L)$
19:$M\leftarrow\text{MaskAugmentation}(M)$
20:$I_{PF}\leftarrow I_{s}\odot M+I_{t}\odot(1-M)$
Figure 2: The overall framework of our proposed method consists of: 1) FIA-USA, which categorizes deepfake artifacts into Face Inconsistency Artifacts (FIA) and Up-Sampling Artifacts (USA). By employing data augmentation techniques, the model is enhanced to concurrently recognize and learn from both categories of artifacts. 2) Region-aware Contrastive Regularization (RCR), which enables the model to focus simultaneously on forged boundaries and spatial artifacts through the contrastive regularization of features from forged and real regions. 3) Automatic Forgery-aware Feature Selection (AFFS), which achieves efficient feature selection by assessing the similarity between the feature maps of each channel in each layer and the mask.
3.4 Loss Function
In addition to FIA-USA, which aims to generate fake training samples from real training samples, we also propose Automatic Forgery aware Feature Selection (AFFS) and Region aware Contrastive Regularization (RCR) based on FIA-USA.
3.4.1 Automatic Forgery-aware Feature Selection
Before the model training begins, AFFS calculates the similarity between the feature map differences of each dimension of the real and fake faces in the each layer of the neural network and the mask $M$, sorts from high to low and selects the top $m_{i}$ dimensions of features of the $i$-th layer to construct a Feature Pyramid Network (FPN) [49]. It is worth noting that AFFS is only performed before the model training starts, and once completed, which dimensions of each layer’s features will be used to construct the FPN is fixed. AFFS enables adaptive feature selection, reduces feature redundancy, enhances feature representativeness and discriminability, and ensures effective Region-aware Contrastive Regularization.
Specifically, given a neural network $\phi$, we define the output of $i$-th layer as $f_{i}=\phi_{i}(I)\in\mathbb{R}^{W_{i}*H_{i}*C_{i}}$ where $I\in\mathbb{R}^{H*W*3}$ represents the input image, $W_{i}$ and $H_{i}$ refer to the size of the feature map, and $C_{i}$ represents the feature dimension of $i$-th layer. Before training begins, we select the top $m_{i}$ channels for the $i$-th layer by minimizing the following loss function.
$\mathcal{L}_{AFFS}^{i,k}=\frac{1}{N}\sum_{n=1}^{N}\|\mathcal{F}(f_{i}^{k}(R_{n})-(f_{i}^{k}(F_{n}))-M_{n}\|_{2}^{2}$ (2)
$\mathcal{F}$ represents normalization and scaling operations, $f_{i}^{k}\in\mathbb{R}^{W_{i}*H_{i}},k\in[0,C_{i}]$, represents the feature map of the $k$-th dimension of the $f_{i}$ , $R_{n}$ and $F_{n}$ are positive and negative samples generated by FIA-USA, $M_{n}$ is the corresponding mask. Finally, we use the features $f_{i}^{\prime}$ selected by AFFS to construct FPN, where $f_{i}^{\prime}\in\mathbb{R}^{W_{i}*H_{i}*m_{i}}$.
3.4.2 Region-aware Contrastive Regularization
Given the network $\phi$ and the Feature Pyramid Network (FPN) $P$, we first define a set of triplets $F$, $R$, $M$, where $F$ and $R\in\mathbb{R}^{H*W*3}$ represents the pseudo-fake face and real face, and $M\in\mathbb{R}^{H*W}$ represents the corresponding mask in FIA-USA used to generate the fake face. We denote the feature maps of real and fake faces as $H_{R}=P(\phi(R))$, $H_{F}=P(\phi(F))$. For $H_{R}$, the areas inside and outside $M$ are real, while for $H_{F}$, the areas inside $M$ are manipulated and the areas outside $M$ are real. We define the region within $M$ as $H_{R}^{in}$, $H_{F}^{in}$ and the region outside $M$ as $H_{R}^{out}$, $H_{F}^{out}$. It is noteworthy that, aligning with LAA-Net [55], we acknowledge that pixels on the blend boundary are influenced by both false pixels within $M$ and real pixels outside $M$, hence we discard pixel features on the blend boundary.
Patch-level Contrastive Regularization. For real faces, the features $f$ of $H_{R}^{in}$ and $H_{R}^{out}$ are consistent, thus we aim to minimize the distance between them. Conversely, for fake faces, the features $f$ of $H_{F}^{in}$ and $H_{F}^{out}$ are distinct, thus we aim to maximize the distance between them. Therefore, the loss function of Patch-level Contrastive Regularization (PCR) can be formulated as:
$\mathcal{L}_{1}=-\log\frac{\sum\limits_{f_{R}\in H_{R}}e^{\delta(f_{R}^{in},f_{R}^{out})/\tau}}{\sum\limits_{f_{R}\in H_{R}}e^{\delta(f_{R}^{in},f_{R}^{out})/\tau}+\sum\limits_{f_{F}\in H_{F}}e^{\delta(f_{F}^{in},f_{F}^{out})/\tau}},$ (3)
where $\delta(\cdot)$ is the normalized cosine similarity between two features, as:
$\delta(f_{1},f_{2})=\sum\limits_{f_{1}^{i}\in f_{1}}\sum\limits_{f_{2}^{i}\in f_{2}}\frac{f_{1}^{i}}{\|f_{1}^{i}\|}\cdot\frac{f_{2}^{i}}{\|f_{2}^{i}\|},$ (4)
where $f^{i}$ represents a pixel feature. $\tau$ is a temperature parameter.
Image-level Contrastive Regularization. Since both the region of real face and fake face outside $M$ are real, we aim to minimize the distance between them and maximize the distance between the fake face and the real face in the area inside $M$. Therefore, the loss function for Image-level Contrastive Regularization (ICR) can be formulated as:
$\mathcal{L}_{2}=-\log\frac{\sum\limits_{f_{out}\in H^{out}}e^{\delta(f_{R}^{out},f_{F}^{out})/\tau}}{\sum\limits_{f^{out}\in H_{out}}e^{\delta(f_{R}^{out},f_{F}^{out})/\tau}+\sum\limits_{f^{in}\in H^{in}}e^{\delta(f_{R}^{in},f_{F}^{in})/\tau}}.$ (5)
Overall Loss. The network is optimized using the following loss:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{BCE}+\lambda_{2}\mathcal{L}_{1}+\lambda_{3}\mathcal{L}_{2},$ (6)
where $\mathcal{L}_{BCE}$ denotes the cross-entropy classification loss. $\mathcal{L}_{BCE},$ $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ are weighted by the hyperparameters $\lambda_{1}$ ,$\lambda_{2}$ and $\lambda_{3}$, respectively.
4 Experiments
4.1 Settings
Detector Backbone Venues CDF-v1 CDF-v2 DFD DFDCP Avg.
FWA [45] Xception CVPRW’18 0.790 0.668 0.740 0.638 0.714
CapsuleNet [56] Capsule ICASSP’19 0.791 0.747 0.684 0.657 0.720
CNN-Aug [30] ResNet CVPR’20 0.742 0.703 0.646 0.617 0.677
Face X-ray [44] HRNet CVPR’20 0.709 0.679 0.766 0.694 0.712
FFD [11] Xception CVPR’20 0.784 0.744 0.802 0.743 0.768
F3Net [60] Xception ECCV’20 0.777 0.798 0.702 0.735 0.749
SPSL [51] Xception CVPR’20 0.815 0.765 0.812 0.741 0.783
SRM [53] Xception CVPR’21 0.793 0.755 0.812 0.741 0.775
CORE [57] Xception CVPRW’22 0.780 0.743 0.802 0.734 0.764
Recce [2] Designed CVPR’22 0.768 0.732 0.812 0.734 0.761
SBI* [66] EfficientNet-B4 CVPR’22 - 0.813 0.774 0.799 -
UCF [84] Xception ICCV’23 0.779 0.753 0.807 0.759 0.774
ED* [1] ResNet-34 AAAI’24 0.818 0.864 - 0.851 -
ProDet† [6] EfficientNet-B4 NeurIPS’24 0.909 0.842 0.848 0.774 0.843
LSDA* [86] EfficientNet-B4 CVPR’24 0.867 0.830 0.880 0.815 0.848
Forensic-Adapter* [10] CLIP (ViT-B/16) CVPR’25 - 0.837 - 0.799 -
Ours EfficientNet-B4 - 0.901 0.867 0.821 0.818 0.852
Table 2: Cross-dataset Evaluations: Cross-dataset evaluations were conducted using the Frame-level AUC. All experiments were trained on the c23 version of FF++ and tested on other datasets. ‘*’ indicates the results are cited from [10] and ‘†’ indicates our reproduction results using the checkpoints provided by the authors, otherwise, the results are from DeepfakeBench [85].
Figure 3: Robustness to Unseen Perturbations. We reported the Video-Level AUC Reduction Ratio for four specific types of perturbations at five different levels.
Datasets.
To evaluate the effectiveness of our proposed method, we conducted extensive experiments on multiple datasets, including both traditional and generative deepfake datasets.
Traditional Deepfake Datasets. FaceForensics++ (FF++) [63], DeepfakeDetection (DFD) [12], Deepfake Detection Challenge (DFDC) [15], preview version of DFDC (DFDCP) [14], and CelebDF (CDF) [47]. FF++ comprises 1,000 original videos and 4,000 fake videos forged by four manipulation methods, namely, Deepfakes(DF) [13], Face2Face(F2F) [72], FaceSwap(FS) [18], and NeuralTextures(NT) [73]. Consistent with the SBI [67], we trained only with real faces from FF++. FF++ offers three levels of compression: raw, lightly compressed (c23), and heavily compressed (c40). Due to constraints regarding storage and download capabilities, we selected the c23 version of FF++, including DFD. Although many previous studies have utilized the same dataset for both training and testing, the preprocessing and experimental configurations can differ, making fair comparisons difficult. Therefore, in addition to testing on the raw data of the aforementioned datasets, we also performed generalization assessments on the unified new benchmark for traditional deepfakes, (DeepfakeBench) [85].
Method Venues uniface e4s facedancer fsgan inswap simswap Avg.
RECCE [2] CVPR 2022 84.2 65.2 78.3 88.4 79.5 73.0 78.1
SBI [67] CVPR 2022 64.4 69.0 44.7 87.9 63.3 56.8 64.4
CORE [57] CVPRW 2022 81.7 63.4 71.7 91.1 79.4 69.3 76.1
IID [34] CVPR 2023 79.5 71.0 79.0 86.4 74.4 64.0 75.7
UCF [84] ICCV 2023 78.7 69.2 80.0 88.1 76.8 64.9 77.5
LSDA [86] CVPR 2024 85.4 68.4 75.9 83.2 81.0 72.7 77.8
CDFA [50] ECCV 2024 76.5 67.4 75.4 84.8 72.0 76.1 75.9
ProgressiveDet [6] NeurIPS 2024 84.5 71.0 73.6 86.5 78.8 77.8 78.7
Ours - 91.8 87.5 83.0 86.3 87.4 91.0 87.8
Table 3: We compared our method with state-of-the-art detection methods on six representative face swapping forgeries in DF40 [87].
Generative Deepfake Datasets.
Diffusion Facial Forgery (DiFF) [5] contains over 500,000 images synthesized using 13 different generation methods under four conditions: Text to Image (T2I), Image to Image (I2I), Face Swapping (FS) and Face Editing (FE). Text to Image encompasses four generation methods: Midjourney [54], Stable Diffusion XL (SDXL) [59], Free-DoM_T [88], and HPS [82]. Image to Image comprises Low Rank Adaptation(LoRA) [32], DreamBooth [64], SDXL Refiner [59], and Free-DoM_I. Face Swapping includes DiffFace [40] and DCFace [40]. Face Editing comprises Imagic [39], Cycle Diffusion(CycleDiff) [80], and Collaborative Diffusion(CoDiff) [35]. Besides, we also conducted experiments on the latest deepfake dataset DF40 [87], which contains the forgery data generated within the FF++ domain, ensuring the evaluation fake methods different but the data domain unchanged.
Evaluation Metrics.
We report the Frame-Level Area Under Curve (AUC) metric on the DeepfakeBench and DIFF dataset, report the Video-Level Area Under Curve (AUC) metric on the traditional deepfake datasets, to compare our proposed method with prior works.
Implementation Details.
We utilize EfficientNetB4 [71] as the default neural network, and employ the SAM [21] optimizer to train it for 50 epochs. The batch size is set to 12. We uniformly sample 32 frames from each video during training. The initial learning rate is 0.001. For data augmentation, in addition to the proposed data augmentation (FIA-USA), RandomHorizontalFlip, RandomCutOut, and AddGaussian Noise were also applied. All experiments were conducted on a single NVIDIA 3090. We also investigate alternative network architectures and their respective outcomes. Empirically, $\lambda_{1},\lambda_{2},\lambda_{3}$ are assigned values of 1, 2.5, 0.25. The temperature parameter $\tau$ is 0.7. We also explored the impact of other variants on the detection results.
4.2 Generalization Evaluation
Our model’s generalization was evaluated on traditional and generation deepfake datasets at both the frame and video levels.
4.2.1 Traditional Deepfake Datasets
Model Venues CDF-v2 DFDC DFDCP
Face X-Ray [44] CVPR’20 - - 0.711
LipForensics [28] CVPR’21 0.824 0.735 -
FTCN [94] ICCV’21 0.869 - 0.740
PCL + I2G [93] ICCV’21 0.900 0.675 0.743
HCIL [25] ECCV’22 0.790 0.692 -
ICT  [17] CVPR’22 0.857 - -
SBI [67] CVPR’22 0.928 0.719 0.855
AltFreezing [79] CVPR’23 0.895 - -
SAM [7] CVPR’24 0.890 - -
LSDA [86] CVPR’24 0.911 0.770 -
CFM [52] TIFS’24 0.897 - 0.802
Ours - 0.941 0.732 0.866
Table 4: Comparison with SOTA methods using the Video-Level AUC. The results of other methods are cited from their original papers.
Frame-level.
We first compare our method with previous work at the frame level. To ensure the fairness of the experiment, we conducted the experiment on the DeepfakeBench, and adhered to the data preprocessing and experimental settings provided by them. For previous work, we utilized the experiment data provided by DeepfakeBench. As shown in Table 2, our method outperforms other methods on the majority of deepfake datasets and competes with state-of-the-art methods on the CDF-v1 dataset.
Video-Level.
In addition to comparing at the Frame-Level, we also compared our method at the Video-Level with state-of-the-art detection algorithms, including various data augmentation methods. Typically, the experimental results are directly taken from the original paper and trained on the c23 version of FF++. FreqBlender [43] did not provide experimental results trained on the FF++ c23 version in their paper, therefore we directly cited experiment results on the FF++ raw version. The comparison results, as shown in Table 4, strongly demonstrate the effective generalization of our method, and our method outperforms other methods, even being comparable to FreqBlender trained on the raw version of FF++.
Method Train Set Test Subset
Real Fake T2I I2I FS FE
Xception [63] 62.43 56.83 85.97 58.64
F3-Net [60] 66.87 67.64 81.01 60.60
EfficientNet [71] 74.12 57.27 82.11 57.20
DIRE [78] 44.22 64.64 84.98 57.72
SBI [67] 80.20 80.40 85.08 68.79
Ours 86.05 84.95 89.42 72.73
Table 5: Comparison with universal deepfake detection methods using the Frame-Level AUC on the DiFF dataset. ‘†’ indicates models that are designed for deepfake detection, while ‘‡’ signifies models intended for general generated image detection.
Setting Test Dataset Avg.
Traditional Generative
CDF-v2 DFDCP FS FE
w FIA-USA w/o Mtm 91.99 87.06 88.00 71.93 84.74
w/o SR 93.25 86.81 80.73 61.07 80.46
w/o AE 92.53 86.37 88.67 70.31 84.47
w RCR w/o PCR 90.12 86.13 89.61 66.87 83.18
w/o ICR 90.85 85.00 89.95 75.01 85.20
w/o RCR 90.14 86.13 89.61 66.87 83.18
w/o FIA-USA 93.18 89.45 80.68 62.99 81.57
w/o AFFS 94.70 86.30 88.33 70.62 84.98
Ours 94.10 86.66 89.42 72.73 85.72
(a) Ablation Study of Method Components. We report the Video-Level AUC for traditional deepfake datasets and the Frame-Level AUC for generative deepfake datasets. Mmr denotes multiple types of masks, SR signifies the process of super-resolution, and AE refers to the reconstruction achieved through autoencoders.
Model Test Dataset Avg.
Traditional Generative
CDF-v2 DFDCP T2I I2I FS FE
Res50 82.45 79.14 74.73 76.94 68.82 58.59 73.44
Res101 85.51 82.35 73.21 74.95 71.26 60.44 74.62
Effb1 91.88 82.30 83.86 85.24 83.73 70.69 82.86
Effb4 94.10 86.66 86.05 84.95 89.42 72.73 85.65
(b) Ablation Study on Model Architectures.
$\lambda_{1},\lambda_{2},\lambda_{3}$ Test Dataset Avg.
Traditional Generative
CDF-v2 DFDCP T2I I2I FS FE
1, 1, 1 90.86 85.01 85.26 84.11 89.91 70.19 84.22
1, 1, 0.25 91.88 86.01 85.26 84.11 89.91 72.19 84.89
1, 2.5, 1 91.82 86.00 85.21 84.08 89.95 71.01 84.67
1, 0.25, 2.5 91.11 85.59 85.09 83.91 89.61 70.92 84.37
1, 2.5, 0.25 94.10 86.66 86.05 84.95 89.42 72.73 85.65
(c) Experiment on Hyperparameter Configuration.
Table 6: Ablation experiment. We conducted ablation experiments on method components, model frameworks, and hyperparameter configuration. The best results are presented in bold.
4.2.2 Generation Deepfake Datasets
As mentioned in Section 2, researchers have found that RGB enhancement methods exhibit limited efficacy when facing generative deepfakes. Here, we compared multiple universal deepfake detection methods on the generative deepfake dataset DIFF, as shown in Table 3 and Table 5. Our method not only exhibits robust generalization capabilities on traditional deepfakes, but also showcases superior performance in generative deepfake detection. Compared to the baseline method, our approach achieved a significant improvement of nearly 5%.
In section G of the appendix, we conducted experiments on more forgery techniques provided by DF40 and DIFF.
4.2.3 Robustness
In Figure 3, we assessed the influence of different levels of random perturbations on the detection performance, the experimental setup adhering to the experimental protocols established in prior studies. We quantified the effects of different levels of Gaussian Blur, Block Wise, Change Contrast, Change Saturation, and JPEG Compression on the detectors. Notably, to gauge accuracy, we employed the reduction rate of Video-Level AUC to assess the robustness of the detector to different degrees of disturbance, which could be expressed as $R_{AUC}=(AUC-AUC_{raw})/AUC_{raw}$, where $AUC_{raw}$ refers to no perturbations. Figure 3 shows that our method is more robust relative to other RGB-based methods. Change Contrast, Change Saturation, and JPEG Compression on the detectors. Notably, to gauge accuracy, we employed the reduction rate of Video-Level AUC to assess the robustness of the detector to different degrees of disturbance, which could be expressed as $R_{AUC}=(AUC-AUC_{raw})/AUC_{raw}$, where $AUC_{raw}$ refers to no perturbations. Figure 3 shows that our method is more robust relative to other RGB-based methods.
Figure 4: GradCAM visualization of fake samples across various deepfake datasets.
4.3 Ablation Study
The impact of method components.
To evaluate the effectiveness of each component in our method, we performed ablation studies on multiple datasets. In Table 6(a), we present the results of our ablation experiments on FIA-USA, Automatic Forgery-aware Feature Selection (AFFS), and Region-aware Contrastive Regularization (RCR), as well as the ablation of components in FIA-USA and RCR. The experimental results in the Table 6(a) underscore the efficacy of each component of our method. Notably, we report Video-Level AUC for traditional deepfake datasets and Frame-Level AUC for generative deepfake datasets.
The impact of model framework. Concurrently, we also examined the influence of various model architectures, as detailed in the Table 6(b). We assessed the performance implications of EfficientNet [71], Resnet [30], and network layers on model performance. To ensure a fair comparison with ResNet, we extended training by an additional 100 epochs to achieve a training loss comparable to that of EfficientNet.
Impact of hyperparameters. We also examined the effect of hyperparameters of the loss function on model performance, as detailed in the Table 6(c). Our analysis of the weights revealed that increasing the weight of $\lambda_{2}$ relative to the weights of cross entropy enhances model performance, whereas decreasing the weight of $\lambda_{3}$ relative to cross entropy also improves performance.
5 Visualizations
We employed GradCAM [95] to identify the regions activated for forgery detection. The visualization results depicted in Figure 4 demonstrate that our model precisely identifies the areas of deepfake forgery. Beyond visualizing tampered traditional deepfake images, we also conducted visualizations on images produced by state-of-the-art generation techniques.
6 Conclusion
In conclusion, this study presents a universal framework for deepfake detection by focusing on common artifacts that span various types of face forgeries. By categorizing deepfake artifacts into Face Inconsistency Artifacts (FIA) and Up-Sampling Artifacts (USA), we enhance the generalization capability of detection models. This targeted approach allows the model to focus on detecting fundamental artifact patterns, and potentially improving its performance across diverse deepfake variations. In doing so, our framework may help address some challenges in current detection methods and could provide a promising foundation for developing more adaptable deepfake detection systems in future research.
Appendix A Appendix
This supplementary material provides:
  • Sec. B: We discussed the classification and basic process of forgery techniques.
  • Sec. C: We reviewed the application of data augmentation methods in forgery detection and discussed more details about FIA-USA.
  • Sec. D: We introduced the construction process of FPN.
  • Sec. E: We provided a detailed introduction to AFFS was provided, along with visual evidence.
  • Sec. F: We provide a visualization example of which boundary pixels were discarded by RCR.
  • Sec. G: We provided test results on more forgery techniques.
  • Sec. H: We provide experimental results combining FIA-USA with other state-of-the-art detection methods.
  • Sec. I: We provided more visual examples.
Appendix B Face forgery techniques
In this section, we discuss the classification of facial forgery technology and the basic process of forgery technology, and at which nodes FIA and USA will be introduced.
B.1 Classfication of facial forgery technology
According to different processing procedures and technical principles, we first divide forgery techniques into two categories:
1). Full Face Synthesis: Generate the entire image/video directly through the image / video generation model, without involving facial cropping and stitching in the basic process. Therefore, the artifacts involved in such forgery techniques mainly include USA and a small amount of FIA, where USA is the dependency between adjacent pixels in the forgery area due to the performance of the generator, and FIA is the inconsistency in the generator’s ability to generate face areas with dense details and background areas with sparse details.
2). Face Swapping. Crop the original face, generate fake regions through a generator (optional), and then stitch it with the target face. Therefore, this forgery technique can be divided into two situations: a. It includes FIA and USA: FIA is mainly caused by cropping and stitching, as well as generator performance, while USA is mainly determined by generator performance. b. It only includes FIA: FIA is mainly caused by ropping and stitching. Furthermore, based on the forged area, we can divide face swapping into two parts: Macro-editing and Micro-editing. Macro-editing can be further classified into two subtypes: Editing based on 4 facial keypoints and Editing based on 81 keypoints. Micro-editing allows for precise adjustments to each facial feature.
Figure 5 provides a detailed overview of our classification concept.
Figure 5: We categorize forgery techniques into two main types: Full face synthesis and Face swapping. Face swapping is further divided into macro-editing (based on 4 or 81 keypoints) and micro-editing.
B.2 Basic Process
In Figure 6, we depict the basic process of each type of forgery technique. We can see from this that the main difference between full face synthesis and face swapping lies in whether they include face cropping and stitching operations, which leads to the emergence of a large number of FIA, and whether the background is generated by the generator. The main difference between macro-editing and micro-editing in face swapping is whether the forged area is the entire face or can finely fabricate facial features. Further, macro editing based on 81 key points is more detailed, and the edges of the forged area are more similar to the edges of a real face, while 4 key points result in the forged area being a square. This discovery is mainly due to a detailed observation of the widely used forgery technique ROPE [31] and roop [65], which provides two types of face cropping techniques.
Figure 6: We illustrates the basic processes of forgery techniques, full face synthesis and face swapping differ in face cropping and stitching, while macro-editing (4 vs 81 keypoints) and micro-editing vary in forgery scope and detail.
B.3 At what stage will FIA and USA be introduced
Figure 7: The operations that led to the occurrence of FIA and USA in the basic process.
In Figure 7, we demonstrate which operations lead to the emergence of FIA and USA in the basic process. The observation of a phenomenon led to the refinement of our FIA-USA design.
Appendix C Details of FIA-USA
C.1 Blend-based data augmentation
In Deepfake detection, a successful approach is to manually construct negative samples for training, where data augmentation in the RGB domain is based on Blend, which can be represented as follows:
$I_{F}=M\odot I_{t}+(1-M)\odot I_{s},$ (7)
Among them, $I_{t}$ represents the target face, $I_{s}$ represents the source face, and $I_{F}$ represents the negative sample. The earliest design [44] was to use facial similarity to find the most suitable $I_{s}$ for $I_{t}$. Later, SBI [67] proposed self-blend to improve the statistical consistency of $I_{F}$. There are also works [68] that separate the background and face of $I_{t}$, add noise to the background, and reconstruct $I_{t}$, using the reconstructed $I_{t}$ for blend. However, to our knowledge, there has been no work so far that considers the introduction of both global and local artifacts, let alone the introduction of FIA at both the full face and regional levels while introducing USA.
C.2 Other types of data augmentation
In recent years, many other data augmentation methods have been proposed, such as frequency domain [43] and latent space [86]. Many studies suggest [43, 86] that compared to data augmentation in other domains, data augmentation in the RGB domain [44, 67, 68] exhibits weak robustness and limited effectiveness. However, we have demonstrated through extensive experiments that our proposed data augmentation in the RGB domain is not weaker or even better than that in other domains. It is worth mentioning that our proposed data augmentation belongs to the image level, and we surprisingly found that it is superior to the most advanced data augmentation methods at the video level [79].
C.3 The detailed process of generating masks in FIA-USA
Figure 8: The generation of macro and micro-editing masks in FIA-USA, using convex hull and erosion / expansion for macro masks and random feature combination for micro masks.
Here we have detailed the processing of generating masks in FIA-USA, as shown in Figure 8. Firstly, we divide mask generation into two types: 1) Macro editing masks and 2) Micro editing masks.
Macro editing mask: We follow the mask generation process in SBI [67], which includes the following steps: a. Calculate the convex hull of 81 facial keypoints to obtain the mask of facial contour b. Erosion or expansion of the mask; Besides, we also considered the case of calculating convex hull based on 4 facial keypoints, which originated from the observation of two methods for extracting raw faces provided by the open-source deepfake project that is truly available [65, 31].
Micro-editing masks: The main method is to simulate the forgery process of manipulating facial details randomly by randomly combining facial features.
C.4 Why use Autoencoder and Super-Resolution models to reconstruct images
Figure 9: We divide the generation models into two categories based on whether facial super-resolution models is needed for post-processing.
Figure 10: We demonstrate through graphical representation that our proposed FIA-USA covers all possible scenarios that may lead to the occurrence of FIA and USA.
As shown in Figure 9, we divide the processing flow of the generative model into two categories: One category generates images of forged regions that are as clear and free of blur as the original image (using Diffusion models and GANs as the main methods, and upsampling the latent code through AE to generate images). And another type (especially older voice driven lip shape modification techniques) will result in blurred tampered areas, requiring additional facial super-resolution models to process the blurred areas. Therefore, we use utoencoder (AE) to reconstruct images or facial Super-Resolution (SR) to process faces, in order to approximate the USA introduced by the generator in the forged area.
C.5 FIA-USA covers all possible situations that may occur in FIA and USA
As shown in the Figure 10, our proposed FIA-USA framework covers every possible situation that may occur in FIA and USA, which is also the most intuitive explanation why our method is effective.
Figure 11: We illustrated the construction process of FPN.
Appendix D Details of FPN
As shown in the Figure 11, our FPN follows the framework proposed by Lin et al., using the output features of the 2nd, 3rd, 4th, and 5th layers of the backbone to construct an FPN, In our experimental setup, the number of feature map channels is 128 and 196, corresponding to EfficientNet [71], and Resnet [30], respectively.
Appendix E Details of AFFS
Here we provide a detailed description of the process of AFFS and a visual explanation of its effectiveness.
E.1 Why is AFFS proposed
Let’s consider a neural network $\phi$ and an FPN $P$, as well as the process of performing RCR on the FPN’s maximum resolution layer. The output of $\phi$ is a set of features $\{d_{1},d_{2},d_{3},d_{4}\}$, $P$ takes this set of features as input to construct a set of feature maps with same dimensions and different resolutions, while RCR performs feature differentiation learning on the feature map with the highest resolution. One intuition is that there is a significant amount of redundancy in the output features of the $\phi$, and not every dimension of the features $d_{i},i\in[1,2,3,4]$ is suitable for constructing FPN for RCR. Therefore, we propose AFFS for feature dimensionality reduction of each feature $d_{i}$ before constructing FPN. Moreover, this feature dimensionality reduction method effectively exploits the potential advantages of our proposed FIA-USA and lays the foundation for the effective execution of RCR.
Figure 12: We provide a visual example to illustrate which pixels will be discarded.
E.2 AFFS
Here, we will provide a more comprehensive explanation of the workflow of AFFS from a different perspective.
1) First, let’s consider multiple positive and negative sample pairs generated by FIA-USA, along with their corresponding masks $\{R_{n},F_{n},M_{n}\}_{n=1}^{N}$. The usual method is to directly construct an FPN using a set of features $f_{i}$ from a neural network $\phi$ as input, where $i$ represents the $i$-th layer of the $\phi$. We can see that this process is not related to FIA-USA, so it cannot fully unleash the potential of FIA-USA, and the constructed feature pyramid also intuitively lays the foundation for RCR.
2) Next, let’s discuss the initialization process of AFFS. For the output of each layer of the neural network, we hope that the difference between the real image and the fake image after normalization is as consistent as possible with the mask $M$ used to represent forged areas. However, we have noticed that not every channel meets this requirement, so we would like to reserve channels that meet the requirements for each layer. Therefore, for $f_{i}\in\mathbb{R}^{W_{i}*H_{i}*C_{i}}$ , we calculate the following loss for each channel on a set of FIA-USA generated samples $\{R_{n},F_{n},M_{n}\}_{n=1}^{N}$ to obtain those excellent dimensions,
$\mathcal{L}_{AFFS}^{i,k}=\frac{1}{N}\sum_{n=1}^{N}\|\mathcal{F}(f_{i}^{k}(R_{n})-(f_{i}^{k}(F_{n}))-M_{n}\|_{2}^{2}$ (8)
$\mathcal{F}$ represents normalization and scaling operations, $f_{i}^{k}\in\mathbb{R}^{W_{i}*H_{i}},k\in[0,C_{i}]$, represents the feature map of the $k-$th dimension of the $f_{i}$ During the training process, we only select the channels reserved for $f_{i}$ during the AFFS initialization process. After AFFS processing, the original features $f_{i}\in\mathbb{R}^{W_{i}*H_{i}*C_{i}}$ will become $f_{i}^{\prime}\in\mathbb{R}^{W_{i}*H_{i}*m_{i}}$
E.3 Visual Explanation of AFFS
Here we provide a set of visual examples of AFFS in Figure 13. The experiment was conducted on the 2-th layer of EfficientNetB4. Unlike the actual situation, in order to enhance readability, we only calculated $L_{AFFS}$ for a single epoch on a single pair of samples. And based on this, sort the 32 dimensions of the features in the 2-th layer. Here, we present the dimensions ranked in the top 14 and bottom 4. In the actual process, we iterate for 5 epochs on all samples on the training set to initialize AFFS. In Table 7, we present the changes in feature dimensions of different backbone features before and after performing AFFS
Figure 13: The result of calculating $L_{AFFS}$ for the 32 dimensions of the 2-th layer features of EfficientNetB4 on a single sample and sorting them by similarity.
Model Input Dimension Output Dimension
Res50 {256,512,1024,2048} {196,384,768,1536}
Res101 {256,512,1024,2048} {196,384,768,1536}
Effb1 {24,40,112,320} {16,24,40,112}
Effb4 {32,56,160,448} {24,32,56,160}
Table 7: Feature dimension parameters of AFFS.
Method Training set CDF DFDCP Avg.
LAA-Net FF++(C23) +SBI 0.840 0.741 0.791
LAA-Net FF++(C23) +FIA-USA 0.875 0.782 0.829
Table 8: We report the Video-level AUC. The combination of our data augmentation method with other Sota detection algorithms.
Appendix F Details of RCR
In the LAA Net [55], a very interesting hypothesis is the pixels on the fake boundary contain both the features of the fake area and the features of the real area, so in our designed RCR, such pixels will be discarded. Figure 12 provides a visual example.
Appendix G Testing on more forgery techniques
In addition to several widely used forgery detection datasets mentioned in the paper, we also conducted experiments on other sota forgery techniques based on the DF40 dataset [87], where real samples were obtained from FF++ and fake samples were constructed using corresponding forgery techniques based on FF++ settings. As shown in the Table 9(a), our proposed forgery detection framework has achieved excellent detection results on the vast majority of forgery methods. Meanwhile, Table 9(b) shows the comparison between our method and other state-of-the-art methods on various forgery techniques in DIFF
Type mobileswap MidJourney faceswap styleclip DiT lia ddim mcnet RDDM StyleGAN3
AUC 0.97 0.97 0.89 0.82 0.66 0.91 0.97 0.84 0.70 0.93
Type e4e pixart whichfaceisreal deepfacelab StyleGANXL heygen facedancer MRAA pirender VQGAN
AUC 0.95 0.93 0.42 0.77 0.41 0.58 0.83 0.85 0.81 0.85
Type StyleGAN2 facevid2vid simswap inswap one_shot_free wav2lip e4s starganv2 tpsm sd2.1
AUC 0.93 0.83 0.91 0.87 0.85 0.76 0.88 0.48 0.83 0.97
Type blendface hyperreenact uniface CollabDiff fsgan danet SiT sadtalker fomm
AUC 0.94 0.80 0.92 0.95 0.86 0.80 0.72 0.74 0.85
(a) We tested all forgery techniques provided on the DF40 dataset and reported Frame-Level AUC.
Method Train Set Test Subset
Real Fake Cycle CoDiff Imagic DiFace DCFace Dream SDXL_R FD_I LoRA Midj FD_T SDXL HPS
F3-Net [60] 36.14 35.00 32.56 25.61 53.29 55.45 65.04 40.90 - - 45.00 61.64 68.96
EfficientNet [71] 56.51 38.38 48.50 64.45 89.13 71.64 65.04 59.93 - - 69.67 64.94 74.63
CNN_Aug [75] 50.31 46.84 75.95 43.10 80.69 58.75 60.10 43.65 - - 47.90 61.95 60.56
SBI [67] 82.32 64.70 49.84 73.64 89.92 75.66 77.73 87.67 89.18 79.56 90.55 80.37 83.30
Ours 84.76 72.61 52.92 76.86 93.63 83.43 80.54 92.68 93.48 82.57 94.85 82.15 88.40
[…]
Table 9: Test results on more forgery techniques.
Appendix H Combining with SoTA detection method
In this section, we plan to combine our proposed FIA-USA data augmentation with other sota image-level detection methods to validate the effectiveness of our proposed data augmentation method. However, due to the limited availability of open-source image level detection methods, we can only provide the performance of our data augmentation method on LAA-Net [55] (CVPR 24) in Table 8. In the original paper, the model was trained on the FF++ dataset in the raw version, while here we trained on the C23 version.
Appendix I More visual examples
From Figure 14 to Figure 16, we demonstrate the fake samples that FIA-USA can generate through different reconstruction methods, including autoencoderr (AE) and facial super-resolution (SR), and multiple types of masks based on three real faces. Figure 17 shows the negative samples randomly generated by FIA-USA during the real training process. Figures 18 to 22 list the images in the DF40 dataset [87] that were misclassified by our method.
Figure 14: More examples of FIA-USA.
Figure 15: More examples of FIA-USA.
Figure 16: More examples of FIA-USA.
Figure 17: More examples of FIA-USA.
Figure 18: Images misclassified by our method in the DF40 dataset.
Figure 19: Images misclassified by our method in the DF40 dataset.
Figure 20: Images misclassified by our method in the DF40 dataset.
Figure 21: Images misclassified by our method in the DF40 dataset.
Figure 22: Images misclassified by our method in the DF40 dataset.